OpenStudy (anonymous):

If you help me, I'll help you....I'm not good at math...please.... Select the best possible first step to solving the following system by first eliminating the y variable. 6x - 5y = -7 10x + 3y = -1 Answer options: Multiply the first equation by 3 and multiply the second equation by 5. Multiply the first equation by 3 and multiply the second equation by -5. Multiply the first equation by -3 and multiply the second equation by 5. Multiply the first equation by 7 and the second equation by -7.

OpenStudy (anonymous):

hey.. how are you?

OpenStudy (anonymous):

Awful

OpenStudy (anonymous):

Multiply the first equation by 3 and multiply the second equation by 5.

OpenStudy (anonymous):

if we do so we have : 18x - 15y = -21 50x + 15y = -5 then if we add the equations we will eliminate y :)

OpenStudy (anonymous):

Thank you, really

OpenStudy (anonymous):

I got B, but I didn't see the -5

OpenStudy (anonymous):

I've been ill, and having to catch up is the worst, especially if you're not good at math

OpenStudy (anonymous):

in fact B is a possibility ! but then you will have to multiply the equation by a negative number which is not nice as multiplying by a positive one and then you will have to subtract the equations and not add them

OpenStudy (anonymous):

so option A is easier

OpenStudy (anonymous):

Thank you love :)

OpenStudy (anonymous):

Do you mind helping me more?

OpenStudy (anonymous):

I'll show you my answers

OpenStudy (anonymous):

sure ill help you

OpenStudy (anonymous):

Thank you:)))

OpenStudy (anonymous):

I'm stuck on this one:

OpenStudy (anonymous):

Choose the system of equations which matches the following graph:

OpenStudy (anonymous):

5x + 2y = -147x + y = 7 5x + 2y = -147x - y = 7 5x - 2y = -147x + y = 7 5x - 2y = -147x - y = 7

OpenStudy (anonymous):

My answer's B

OpenStudy (anonymous):

both graphs have the point (0,-7)

OpenStudy (anonymous):

so we can eliminate the first one since plugging 0,-7 gives false statement : 7x + y = 7 -7 =7 ..

OpenStudy (anonymous):

we can eliminate the third: 5x - 2y = -14 7x + y = 7 14 = -14 and the fourth : 5x - 2y = -14 7x - y = 7 14 = -14

OpenStudy (anonymous):

so yes.. answer is B :)

OpenStudy (anonymous):

Yes, I have that figured out, except where I'm confused at is whether we'd be adding or subtracting in the firs portion

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!!!:)))))))))))))))))))))))))))))

OpenStudy (anonymous):

I got it right!!!!!!!

OpenStudy (anonymous):

I have another one to check. it's inequalities...which I'm disturbingly horrible at

OpenStudy (anonymous):

Choose the system of inequalities that matches the graph below:

OpenStudy (anonymous):